OpenStudy (anonymous):

the total curb weight of a truck, a car a, and a car B is 11,250 lb. The weight of car A is 15 lb less than the weight of two B cars. The truck weighs 2145 lb more than car A. What is the curb weight of each vehicle? HELP ME PLEASE

OpenStudy (anonymous):

Let the weight of car A = x Let the weight of car B = y Let the weight of truck = z x+y + z = 11,250 x-15=2y z+2145=x

OpenStudy (anonymous):

Now solve. I have to go, Good luck.

OpenStudy (anonymous):

thank you baby!!